# Elliptic asymptotic representation of the fifth Painlevé transcendents (corrected version)

### v8

@@ 5.3. Expression of $B_{\phi}(t)$
The justification scheme is described as follows. Let $\mathcal{M}_{0}\in\mathbf{M}^{*}$ be a given monodromy data, and $K(\varepsilon_{0}):$ $\|\mathcal{M}-\mathcal{M}_{0}\|\leq\varepsilon_{0}$ a given compact ball with centre at $\mathcal{M}_{0}$. By the property above combined with the compactness there exists positive numbers $T_{\infty}$ and $C_{0}$ such that, for every $\mathcal{M}\in K(\varepsilon_{0})$, $\|\mathcal{M}-(\mathcal{M})_{\mathrm{as}}\|\leq C_{0}|t|^{-\delta}$ if $|t|\geq T_{\infty}.$ Note that $f(t,\mathcal{M}):=\mathcal{M}_{0}-(\mathcal{M})_{\mathrm{as}}(t)+\mathcal{M}$ is a map $f$: $K(\varepsilon_{0})\to K(\varepsilon_{0})$ continuous in $\mathcal{M}\in K(\varepsilon_{0})$ if $\|\mathcal{M}-(\mathcal{M})_{\mathrm{as}}(t)\|\leq\varepsilon_{0},$ i.e. if $|t|^{\delta}\geq C_{0}/\varepsilon_{0}$. Then by Brouwer’s fixed point theorem, for each $|t|\geq\max\{T_{\infty},(C_{0}/\varepsilon_{0})^{1/\delta}\}$, there exists a fixed point $\mathcal{M}_{*}=\mathcal{M}_{*}(t)\in K(\varepsilon_{0})$ such that $f(t,\mathcal{M}_{*})=\mathcal{M}_{*}$ i.e. $(\mathcal{M}_{*})_{\mathrm{as}}(t)=\mathcal{M}_{0}$. (As will be shown later $\mathcal{M}_{*}$ is a unique fixed point.) This implies that $\|\mathcal{M}_{*}-\mathcal{M}_{0}\|=\|\mathcal{M}_{*}-(\mathcal{M}_{*})_{\mathrm{as}}(t)\|\leq C_{0}|t|^{-\delta}$ and hence $\mathcal{M}_{*}=\mathcal{M}_{0}+O(t^{-\delta})$ for $|t|\geq\max\{T_{\infty},(C_{0}/\varepsilon_{0})^{1/\delta}\}$. Then $(\mathcal{M}_{*})_{\mathrm{as}}(t)=\mathcal{M}_{0}$ leads us to
@@
which realises isomonodromy deformation with the invariant monodromy data $\mathcal{M}_{0}$, and then $y^{*}_{\mathrm{as}}=(d/dt)y_{\mathrm{as}}$. This provides the desired expression $y_{\mathrm{as}}(\mathcal{M}_{0},t+O(t^{-\delta}))$ in our main result. It remains to show the uniqueness of $\mathcal{M}_{*}.$ To do so suppose that there exist $\mathcal{M}_{*1}$ and $\mathcal{M}_{*2}$ such that $\mathcal{M}_{0}=(\mathcal{M}_{*1})_{\mathrm{as}}(t)=(\mathcal{M}_{*2})_{\mathrm{as}}(t),$ and set $y_{1}=y_{\mathrm{as}}(\mathcal{M}_{*1},t),$ $y_{2}=y_{\mathrm{as}}(\mathcal{M}_{*2},t)$. Let $Y_{1}(\lambda)$ and $Y_{2}(\lambda)$ be matrix solutions solving (3.1) with $\mathcal{B}(t,\lambda)$ containing $(y_{1},y_{1}^{\prime})$ and $(y_{2},y_{2}^{\prime})$, respectively. Since $Y_{1}(\lambda)$ and $Y_{2}(\lambda)$ yield the same monodromy, $Y_{1}(\lambda)^{-1}Y_{2}(\lambda)\equiv\Gamma_{0}$ for some $\Gamma_{0}\in SL_{2}(\mathbb{C})$, and hence $(dY_{1}/d\lambda)Y_{1}^{-1}=(dY_{2}/d\lambda)Y_{2}^{-1}$, implying $y_{1}=y_{2}$, which contradicts the supposition.
@@
Suppose that $0<|\phi-\pi|<\pi/2$, i.e. $\pi/2<\phi<3\pi/2.$ Recall that $\mu(\lambda)$ is on the Riemann surface $\mathbb{P}_{+}\cup\mathbb{P}_{-}$ glued along the cuts $\lambda_{1}^{0},\lambda_{1}]$ and $\lambda_{2},\lambda_{2}^{0}]$. Note that $A_{\phi}=A_{\phi-\pi}$ by Lemma 8.14. Let $\mathcal{S}(\pi/2,3\pi/2)$ on $\mathbb{P}_{+}^{\infty}$ be the limit Stokes graph as described in Figure 6.1 (a), (b), in which $\hat{\mathbf{c}}^{\infty}_{3}$ joins $\lambda_{1}$ or $\lambda_{2}$ to $e^{5\pi i/2}\infty$, and $\mathbf{c}^{\infty}_{3}$ joins $\lambda_{1}$ or $\lambda_{2}$ to $e^{3\pi i/2}\infty$. The anticlockwise $\pi$-radian rotation of the Stokes graph $\mathcal{S}(-\pi/2,\pi/2)$ for $0<|\phi|<\pi/2$ as in Figure 4.1 results in $\mathcal{S}(\pi/2,3\pi/2)$. The curve $\mathbf{c}^{\infty}_{3}$ corresponds to $\mathbf{c}^{\infty}_{1}$ or $\mathbf{c}^{\infty}_{2}$. Let the loops $\breve{l}_{0}$, $\breve{l}_{1}$ be the results of the same rotation of $\hat{l}_{0}$, $\hat{l}_{1}$ in Figure 3.1. The loops $\breve{l}_{0}$, $\breve{l}_{1}$ and the starting point $\breve{p}_{\mathrm{st}}$ are as in Figure 6.1 (c), and $\arg(\breve{p}_{\mathrm{st}})=3\pi/2$.
@@
$\hat{J}_{3}=\lim_{\begin{subarray}{c}\lambda\to\infty\\ \lambda\in\hat{\mathbf{c}}^{\infty}_{3}\end{subarray}}\biggl(\int^{\lambda}_{\lambda_{2}}\Lambda_{3}(\tau)d\tau-\tfrac{1}{4}(t\lambda-2\theta_{\infty}\log\lambda)\sigma_{3}\biggr),$
@@
the first of which immediately follows from $\Gamma_{3}\breve{M}^{0}=S^{-1}_{3}$. Thus the conclusion of the theorem follows.
For general angles $\phi$ such that $|\phi-k\pi|<\pi/2$ $(k\in\mathbb{Z})$, denote by $\hat{l}^{(k)}_{0}$, $\hat{l}^{(k)}_{1}$ and $\mathcal{S}(k\pi-\pi/2,k\pi+\pi/2)$ the $k\pi$-rotation of $\hat{l}_{0},$ $\hat{l}_{1}$ and $\mathcal{S}(-\pi/2,\pi/2)$, respectively. Let $\tilde{Y}^{p}(t,\lambda)$ be the matrix solution of (3.4) admitting the same asymptotic representation as (3.6) in the sector $|\arg\lambda-2p\pi-\pi/2|<\pi,$ and let $M^{0}_{p}$ and $M^{1}_{p}$ be the monodromy matrices given by the analytic continuations of $\tilde{Y}^{p}(t,\lambda)$ along $\hat{l}^{(2p)}_{0}$ and $\hat{l}^{(2p)}_{1}$, respectively. Especially, $\tilde{Y}^{0}(t,\lambda)=Y(t,\lambda),$ $\hat{l}^{(0)}_{0}=\hat{l}_{0},$ $\hat{l}^{(0)}_{1}=\hat{l}_{1},$ $M^{0}_{0}=M^{0},$ $M^{1}_{0}=M^{1}.$ Then $M^{0}_{p}$ and $M^{1}_{p}$ are represented as in Remark 2.5, and using $M^{0}_{p},$ $M^{1}_{p}$ we may reduce the general case to the one $0<|\phi|<\pi/2$, to which Theorem 2.1 applies. Let $\breve{Y}^{p}(t,\lambda)$ be the matrix solution in the sector $|\arg\lambda-(2p+1)\pi-\pi/2|<\pi$ such that $\breve{Y}^{0}(t,\lambda)=Y_{3}(t,\lambda)$. The analytic continuations of $\breve{Y}^{p}(t,\lambda)$ along $\hat{l}^{(2p+1)}_{0}$ and $\hat{l}^{(2p+1)}_{1}$ yield $\breve{M}^{0}_{p}$ and $\breve{M}^{1}_{p}$ as in Remark 2.5, which lead to the case $0<|\phi-\pi|<\pi/2$ of Theorem 2.2 (cf. Remark 2.5).
@@ 7. System equivalent to $(\mathrm{P}_{\mathrm{V}})$ and another approach to $B_{\phi}(t)$
(7.2) $L=L(x):=$ $\frac{(y^{\prime})^{2}}{y(y-1)^{2}}-\frac{y}{(y-1)^{2}}+2(1-\theta_{0}-\theta_{1})\frac{x^{-1}}{y-1}$
@@
In (7.2) set
$\psi=\frac{y+1}{y-1},\quad x=e^{i\phi}t$
(now we return to the variable $x$). The function $L=L(x)$ given by (7.2) is written in terms of $\psi$:
@@
and $b_{0}(e^{i\phi}t)-e^{i\phi}B_{\phi}(t)\ll t^{-2/9+\varepsilon}$ in $S(\phi,t_{\infty},\kappa_{0},\delta_{0})$.
@@ 8. Modulus $A_{\phi}$ and the Boutroux equations
we have $\mathrm{Im\,}(1/\mathcal{I}(A))=-(2\pi)^{-1}\log|A|\,\,(1+o(1))$ as $A\to\infty,$ which implies
@@
Since $I_{\mathbf{a}}(0)=0,$ $I_{\mathbf{b}}(0)=2i,$ $A=0$ solves (BE)ϕ=0. Conversely we may give the uniqueness lemma, which is crucial in discussing (BE)ϕ. This is proved by an argument similar to that in [27, §7]).
@@
$(I_{0}^{\beta}+\tilde{I}^{0}_{\beta})+(\tilde{J}^{-\beta}_{0}+J^{0}_{-\beta})=\frac{2i}{A-\overline{A}}\int^{\beta}_{0}\biggl(\sqrt{\frac{A-(\alpha+it)^{2}}{1-(\alpha+it)^{2}}}-\overline{\sqrt{\frac{A-(\alpha+it)^{2}}{1-(\alpha+it)^{2}}}}\biggr)dt\in i\mathbb{R},$
@@
Since $I_{\mathbf{a}}(1)=4,$ $I_{\mathbf{b}}(1)=0,$ the number $A=1$ solves (BE)ϕ=±π/2. Observe that $\mathrm{Re\,}iI_{\mathbf{b}}(A)=0$ implies $I_{\mathbf{b}}(A)=-I_{\mathbf{b}}(\overline{A}).$ Then, similarly we have
@@
By the uniqueness above we easily have
@@
By Lemmas 8.9, 8.11, 8.15, Proposition 8.13 and Remark 8.1, we have
@@
By the properties above the trajectory of $\Gamma_{0}$ of $A_{\phi}$ for $|\phi|\leq\pi/2$ is as in Figure 8.2, (a).
@@
Figure 8.2. Trajectories of $A_{\phi}$ and conjectural $A^{1/2}_{\phi}$
Thus we have
@@
It is easily verified that $0<\mathrm{Re\,}A^{1/2}_{\phi}<1$ for $0<|\phi|<\pi/2,$ if the trajectory $\Gamma_{0}=\{A_{\phi};\,|\phi|\leq\pi/2\}$ is contained in $P_{0}=\{x+iy;\,y^{2}\leq 4(1-x)\}.$ For some $\varepsilon_{1}\geq\varepsilon_{0}>0,$ by Lemma 8.10, the set $P_{0}$ contains the arc $\{A_{\phi};\,|\phi-\pi/2|<\varepsilon_{0}\}$, and $\{A_{\phi};\,|\phi|<\pi/2-\varepsilon_{1}\}$ as well. It is likely that the trajectory of $A^{1/2}_{\phi}$ is as in Figure 8.2, (b). It remains, however, the possibility of the existence of $A_{\phi}\in\Gamma_{0}$ such that $\mathrm{Re\,}A_{\phi}^{1/2}>1$ and $|\mathrm{Im\,}A^{1/2}_{\phi}|>\varepsilon_{2}>0.$ To deny this possibility it is necessary to show that, say, $\mathrm{Im\,}\mathcal{I}(A)\not=0$ on $\{A;\mathrm{Re\,}A^{1/2}=1\}\setminus\{A=1\}=\partial P_{0}\setminus\{(1,0)\}.$

### v9

@@ 5.3. Expression of $B_{\phi}(t)$
The justification scheme is described as follows. Let $\mathcal{M}_{0}$ be a given monodromy data, and $K(\varepsilon_{0}):$ $\|\mathcal{M}-\mathcal{M}_{0}\|\leq\varepsilon_{0}$ a given compact ball with centre at $\mathcal{M}_{0}$. By the property above combined with the compactness, there exists positive numbers $T_{\infty}$ and $C_{0}$ such that, for every $\mathcal{M}\in K(\varepsilon_{0})$, $\|\mathcal{M}-(\mathcal{M})_{\mathrm{as}}\|\leq C_{0}|t|^{-\delta}$ if $|t|\geq T_{\infty}.$ Note that $f(t,\mathcal{M}):=\mathcal{M}_{0}-(\mathcal{M})_{\mathrm{as}}(t)+\mathcal{M}$ is a map $f$: $K(\varepsilon_{0})\to K(\varepsilon_{0})$ continuous in $\mathcal{M}\in K(\varepsilon_{0})$ if $\|\mathcal{M}-(\mathcal{M})_{\mathrm{as}}(t)\|\leq\varepsilon_{0},$ i.e. if $|t|^{\delta}\geq C_{0}/\varepsilon_{0}$. Then by Brouwer’s fixed point theorem, for each $|t|\geq\max\{T_{\infty},(C_{0}/\varepsilon_{0})^{1/\delta}\}$, there exists a fixed point $\mathcal{M}_{*}=\mathcal{M}_{*}(t)\in K(\varepsilon_{0})$ such that $f(t,\mathcal{M}_{*})=\mathcal{M}_{*}$ i.e. $(\mathcal{M}_{*})_{\mathrm{as}}(t)=\mathcal{M}_{0}$. (As will be shown later $\mathcal{M}_{*}$ is a unique fixed point.) This implies that $\|\mathcal{M}_{*}-\mathcal{M}_{0}\|=\|\mathcal{M}_{*}-(\mathcal{M}_{*})_{\mathrm{as}}(t)\|\leq C_{0}|t|^{-\delta}$ and hence $\mathcal{M}_{*}=\mathcal{M}_{0}+O(t^{-\delta})$ for $|t|\geq\max\{T_{\infty},(C_{0}/\varepsilon_{0})^{1/\delta}\}$. Then $(\mathcal{M}_{*})_{\mathrm{as}}(t)=\mathcal{M}_{0}$ leads us to
@@
which realises isomonodromy deformation with the invariant monodromy data $\mathcal{M}_{0}$, and then $y^{*}_{\mathrm{as}}=(d/dt)y_{\mathrm{as}}$. This provides the desired expression $y_{\mathrm{as}}(\mathcal{M}_{0},t+O(t^{-\delta}))$ in our main result. It remains to show the uniqueness of $\mathcal{M}_{*}.$ To do so suppose that there exist $\mathcal{M}_{*1}$ and $\mathcal{M}_{*2}$ such that $\mathcal{M}_{0}=(\mathcal{M}_{*1})_{\mathrm{as}}(t)=(\mathcal{M}_{*2})_{\mathrm{as}}(t),$ and set $y_{1}=y_{\mathrm{as}}(\mathcal{M}_{*1},t),$ $y_{2}=y_{\mathrm{as}}(\mathcal{M}_{*2},t)$. Let $V_{1}(\lambda)$ and $V_{2}(\lambda)$ be matrix solutions solving (3.4) with $\mathcal{B}(t,\lambda)$ containing $(y_{1},y_{1}^{\prime})$ and $(y_{2},y_{2}^{\prime})$, respectively. Since $V_{1}(\lambda)$ and $V_{2}(\lambda)$ yield the same monodromy, by the uniqueness of Proposition 3.2 we have $(dV_{1}/d\lambda)V_{1}^{-1}=(dY^{(2)}/d\lambda)V_{2}^{-1}$, implying $y_{1}=y_{2}$, which contradicts the supposition.
@@
Suppose that $0<|\phi-\pi|<\pi/2$, i.e. $\pi/2<\phi<3\pi/2.$ Recall that $\mu(\lambda)$ is on the Riemann surface $\mathbb{P}_{+}\cup\mathbb{P}_{-}$ glued along the cuts $\lambda_{1}^{0},\lambda_{1}]$ and $\lambda_{2},\lambda_{2}^{0}]$. Note that $A_{\phi}=A_{\phi-\pi}$ by Lemma 8.14. Let $\mathcal{S}(\pi/2,3\pi/2)$ on $\mathbb{P}_{+}^{\infty}$ be the limit Stokes graph as described in Figure 6.1 (a), (b), in which $\hat{\mathbf{c}}^{\infty}_{3}$ joins $\lambda_{1}$ or $\lambda_{2}$ to $e^{5\pi i/2}\infty$, and $\mathbf{c}^{\infty}_{3}$ joins $\lambda_{2}$ or $\lambda_{1}$ to $e^{3\pi i/2}\infty$. The anticlockwise $\pi$-radian rotation of the Stokes graph $\mathcal{S}(-\pi/2,\pi/2)$ for $0<|\phi|<\pi/2$ as in Figure 4.1 results in $\mathcal{S}(\pi/2,3\pi/2)$. The curve $\mathbf{c}^{\infty}_{3}$ corresponds to $\mathbf{c}^{\infty}_{1}$ or $\mathbf{c}^{\infty}_{2}$. Let the loops $\breve{l}_{0}$, $\breve{l}_{1}$ be the results of the same rotation of $\hat{l}_{0}$, $\hat{l}_{1}$ in Figure 3.1. The loops $\breve{l}_{0}$, $\breve{l}_{1}$ and the starting point $\breve{p}_{\mathrm{st}}$ are as in Figure 6.1 (c), and $\arg(\breve{p}_{\mathrm{st}})=3\pi/2$.
@@
$\hat{J}_{3}=\lim_{\begin{subarray}{c}\lambda\to\infty\\ \lambda\in\hat{\mathbf{c}}^{\infty}_{3}\end{subarray}}\biggl(\int^{\lambda}_{\lambda_{2}}\Lambda_{3}(\tau)d\tau-\tfrac{1}{4}(t\lambda-2\theta_{\infty}\log\lambda)\sigma_{3}\biggr).$
@@
the first of which immediately follows from $\Gamma_{3}\breve{M}^{0}=S^{-1}_{3}$. From these relations the phase shift $\breve{x}_{0}$ is derived by a procedure analogous to that for $x_{0}$. Thus Theorem 2.2 is obtained.
For $\phi$ such that $|\phi-k\pi|<\pi/2$ $(k\in\mathbb{Z})$, denote by $\hat{l}^{(k)}_{0}$, $\hat{l}^{(k)}_{1}$ and $\mathcal{S}(k\pi-\pi/2,k\pi+\pi/2)$ the $k\pi$-rotation of $\hat{l}_{0},$ $\hat{l}_{1}$ and $\mathcal{S}(-\pi/2,\pi/2)$, respectively. Let $\tilde{Y}^{p}(t,\lambda)$ be the canonical solution of (3.4) admitting the same form as of (3.6) in the sector $|\arg\lambda-2p\pi-\pi/2|<\pi,$ and let $M^{0}_{p}$ and $M^{1}_{p}$ be the monodromy matrices given by the analytic continuations of $\tilde{Y}^{p}(t,\lambda)$ along $\hat{l}^{(2p)}_{0}$ and $\hat{l}^{(2p)}_{1}$, respectively. Especially, $\tilde{Y}^{0}(t,\lambda)=Y(t,\lambda),$ $\hat{l}^{(0)}_{0}=\hat{l}_{0},$ $\hat{l}^{(0)}_{1}=\hat{l}_{1},$ $M^{0}_{0}=M^{0},$ $M^{1}_{0}=M^{1}.$ Then $M^{0}_{p}$ and $M^{1}_{p}$ are as in Remark 2.5, and $M^{0}_{p},$ $M^{1}_{p}$ are calculated on $\mathcal{S}(2p\pi-\pi/2,2p\pi+\pi/2)$. Thus this is reduced to the situation, to which Theorem 2.1 applies. For the canonical solution $\breve{Y}^{p}(t,\lambda)$ in the sector $|\arg\lambda-(2p+1)\pi-\pi/2|<\pi$ such that $\breve{Y}^{0}(t,\lambda)=Y_{3}(t,\lambda)$, the analytic continuations along $\hat{l}^{(2p+1)}_{0}$ and $\hat{l}^{(2p+1)}_{1}$ yield $\breve{M}^{0}_{p}$ and $\breve{M}^{1}_{p}$ as in Remark 2.5. Calculation of $\breve{M}^{0}_{p}$ and $\breve{M}^{1}_{p}$ on $\mathcal{S}((2p+1)\pi-\pi/2,(2p+1)\pi+\pi/2)$ leads to the results corresponding to Theorem 2.2 (cf. Remark 2.5).
@@ 7. System equivalent to $(\mathrm{P}_{\mathrm{V}})$ and another approach to $B_{\phi}(t)$
$L=L(x):=$ $\frac{(y^{\prime})^{2}}{y(y-1)^{2}}-\frac{y}{(y-1)^{2}}+2(1-\theta_{0}-\theta_{1})\frac{x^{-1}}{y-1}$
@@
Furthermore $L=L(x)$ is written in terms of $\psi:=(y+1)/(y-1)$:
@@
and $b_{0}(e^{i\phi}t)-e^{i\phi}B_{\phi}(t)\ll t^{-\delta}$ in $S(\phi,t_{\infty},\kappa_{0},\delta_{0})$.
@@ 8. Modulus $A_{\phi}$ and the Boutroux equations
we have $\mathrm{Im\,}(1/\mathcal{I}(A))=-(2\pi)^{-1}\log|A|\,\,(1+o(1))$ as $A\to\infty.$
@@
Since $I_{\mathbf{a}}(0)=0$ and $I_{\mathbf{b}}(0)=2i,$ $A=0$ solves (BE)ϕ=0. Conversely we may give the uniqueness lemma, which is crucial in discussing (BE)ϕ. This is proved by an argument similar to that in [27, §7].
@@
$(I_{0}^{\beta}+\tilde{I}^{0}_{\beta})+(\tilde{J}^{-\beta}_{0}+J^{0}_{-\beta})=\frac{2i}{A-\overline{A}}\int^{\beta}_{0}\biggl(\sqrt{\frac{A-(\alpha+it)^{2}}{1-(\alpha+it)^{2}}}-\overline{\sqrt{\frac{A-(\alpha+it)^{2}}{1-(\alpha+it)^{2}}}}\biggr)dt\in i\mathbb{R}$
@@
Since $I_{\mathbf{a}}(1)=4,$ $I_{\mathbf{b}}(1)=0,$ the number $A=1$ solves (BE)ϕ=±π/2. Observe that $\mathrm{Re\,}iI_{\mathbf{b}}(A)=0$ implies $I_{\mathbf{b}}(A)=-I_{\mathbf{b}}(\overline{A}).$ Then, similarly we have the following.
@@
By the uniqueness above we easily have the following.
@@
Lemmas 8.9, 8.11, 8.15, Proposition 8.13 and Remark 8.1 leads to the following.
@@
By the properties above the trajectory of $A_{\phi}$ for $|\phi|\leq\pi/2$ is as in Figure 8.2, (a).
@@
Figure 8.2. Rough drawings of the trajectories $A_{\phi}$ and $A^{1/2}_{\phi}$
The properties of $A_{\phi}$ are summarised as follows.
@@
By Lemma 8.9, around $\phi=0,$ $A_{\phi}=x(\phi)+iy(\phi)$ is expressed as $A_{\phi}=\pm i2^{3/2}(xL(x)^{-1})^{1/2}(1+o(1))$ and $A_{\phi}^{1/2}=\pm e^{\pi i/4}2^{3/4}(xL(x)^{-1})^{1/4}(1+o(1))$ as $x=x(\phi)=4\phi^{2}L(\phi)^{-1}(1+o(1))\to 0$, where $L(x)=|\log x|.$ By Lemma 8.10, around $\phi=\pm\pi/2$, $A_{\phi}=1-\frac{1}{4}y^{2}L(y)(1+o(1))+iy$ and $A_{\phi}^{1/2}=1-\frac{1}{8}y^{2}L(y)(1+o(1))+iy/2$ as $y=y({\phi})=-4\tilde{\phi}L(\tilde{\phi})^{-1}(1+o(1))\to 0$ $(\tilde{\phi}=\phi\mp\pi/2)$. Taking these local shapes around $\phi=0,$ and $\pm\pi/2$ into account, which are important in finding the Stokes graph, we have a rough drawing of the trajectory $A_{\phi}^{1/2}$ as in Figure 8.2, (b).